, Question 3 Not yet answered Points out of 4.00 Flag question compute the standard deviation σ or ages of British nurses in 1851 Assume that the table below shows the age distribution of nurses in Great Britain in 1851 your answer to nearest hundredth. Round ge range (yr) 20-29 0-69 70-79 75.5 9.0% dpoint (n) 64.5 of nures 8% 8% 9.6% 1% 9% 8% Select one: a. 3.27 b. 13.69 C. 53.76 d. 17.25 е. 1.53
Answer #1
x P(x) xP(x) x2P(x)
24.5 0.058 1.421 34.815
34.5 0.098 3.381 116.645
44.5 0.196 8.722 388.129
54.5 0.291 15.860 864.343
64.5 0.249 16.061 1035.902
74.5 0.09 6.705 499.523
84.5 0.018 1.521 128.525
total 53.670 3067.880
E(x) =μ= ΣxP(x) = 53.6700
E(x2) = Σx2P(x) = 3067.8800
Var(x)=σ2 = E(x2)-(E(x))2= 187.411
std deviation=         σ= √σ2 = 13.69

option b is correct